OpenStudy (anonymous):

If 37.4 grams of water decomposes at 297 Kelvin and 1.30 atmospheres, how many liters of oxygen gas can be produced? Show all of the work used to solve this problem.

OpenStudy (iheart.jesus):

2 H2O = 2 H2 + O2 moles H2O = 37.4 g / 18.02 g/mol=2.08 the ratio between H2O and O2 is 2 : 1 moles O2 produced = 2.08/ 2 = 1.04 V = nRT/p = 1.04 x 0.08206 x 297 / 1.30=19.5 L I really hope this helps.

OpenStudy (anonymous):

19.5